# Half canonical rings of Gorenstein spin curves of genus two

### v1

@@ 5.D.3. Type $B(2)$
$5$ $\langle x_{1}^{5},x_{1}v\rangle\oplus\langle x_{2}^{5},x_{2}v\rangle$
$6$ $\langle x_{1}^{6},x_{1}^{2}v\rangle\oplus\langle x_{2}^{6},x_{2}^{2}v\rangle$ $\langle u\rangle$
$7$ $\langle x_{1}^{7},x_{1}^{3}v\rangle\oplus\langle x_{2}^{7},x_{2}^{3}v\rangle$ $\langle x_{1}u\rangle\oplus\langle x_{2}u\rangle$
$8$ $\langle x_{1}^{8},x_{1}^{4}v\rangle\oplus\langle x_{2}^{8},x_{2}^{4}v\rangle\oplus\langle v^{2}\rangle$ $\langle x_{1}^{2}u\rangle\oplus\langle x_{2}^{2}u\rangle$
$9$ $\langle x_{1}^{9},x_{1}^{5}v,x_{1}v^{2}\rangle\oplus\langle x_{2}^{9},x_{2}^{5}v,x_{2}v^{2}\rangle$ $\langle x_{1}^{3}u\rangle\oplus\langle x_{2}^{3}u\rangle$
@@ 5.E.1. $A(0)$
The set of points in $\mathbb{P}$ where $\mathcal{O}_{\mathbb{P}}(2)$ is not invertible is $\{y_{1}=y_{2}=0\}$ and by our assumptions it does not intersect $C$. Since
@@
and we have the resolution
$0\to{\mathcal{O}}_{\mathbb{P}}(-12)\to{\mathcal{O}}_{\mathbb{P}}(-6)^{\oplus 2}\to{\mathcal{O}}_{\mathbb{P}}\to{\mathcal{O}}_{C}\to 0$
we see that $\omega_{C}\cong\mathcal{O}_{{\mathbb{P}}}(2){\raisebox{-0.1pt}{$|_{C}$}}$ is invertible at all points of $C$, namely $C$ is Gorenstein. Finally, $h^{0}({\mathcal{L}})=h^{0}({\mathcal{O}}_{{\mathbb{P}}}(1))=0$.
@@ 5.E.3. $B(0)$
Computing the resolution of this ideal (e.g., taking the Eagon–Northcott complex or using MAGMA or Macaulay2), one sees that our condition on $g_{8}$ implies that it is a prime ideal, and computing the Hilbert series we see that it defines an irreducible curve $D_{1}$ with $p_{a}(D_{1})=1$. In particular $D_{1}$ is Gorenstein by [Cat82, Lemma 1.19]. An analogous computation for $C_{2}$ shows that $C$ contains $D_{1}\cup D_{2}$, where $D_{1}$ and $D_{2}$ are irreducible Gorenstein curves with arithmetic genus 1 and $C=D_{1}\cup D_{2}$ except possibly at the intersection of $C$ with $\{y_{1}=y_{2}=0\}$. This intersection consists only of the point $P:=[0,0,0,0,1,0,0,\xi]$, where $\alpha\xi^{3}=1$, and an explicit computation shows that $C$ has a node at $P$. Hence $C$ is reduced at all points, $C_{i}=D_{i}$, $i=1,2$, and $C=C_{1}\cup C_{2}$. Summing up, we have shown that $C$ is a reduced Gorenstein curve of genus two of type $B$. Hence $\omega_{C}$ is a line bundle.

### v2

@@ 5.D.3. Type $B(2)$
55 $\langle x_{1}^{5},x_{1}v\rangle\oplus\langle x_{2}^{5},x_{2}v\rangle$
66 $\langle x_{1}^{6},x_{1}^{2}v\rangle\oplus\langle x_{2}^{6},x_{2}^{2}v\rangle$ $\langle u\rangle$
77 $\langle x_{1}^{7},x_{1}^{3}v\rangle\oplus\langle x_{2}^{7},x_{2}^{3}v\rangle$ $\langle x_{1}u\rangle\oplus\langle x_{2}u\rangle$
88 $\langle x_{1}^{8},x_{1}^{4}v\rangle\oplus\langle x_{2}^{8},x_{2}^{4}v\rangle\oplus\langle v^{2}\rangle$ $\langle x_{1}^{2}u\rangle\oplus\langle x_{2}^{2}u\rangle$
99 $\langle x_{1}^{9},x_{1}^{5}v,x_{1}v^{2}\rangle\oplus\langle x_{2}^{9},x_{2}^{5}v,x_{2}v^{2}\rangle$ $\langle x_{1}^{3}u\rangle\oplus\langle x_{2}^{3}u\rangle$
@@ 5.E.1. $A(0)$
The set of points in $\mathbb{P}$ where $\mathcal{O}_{\mathbb{P}}(2)$ is not invertible is $\{y_{1}=y_{2}=0\}$ and by our assumptions it does not intersect $C$. By the very definition of the ideal $I$, which is generated by two polynomials of degree 6 having a relation in degree 12, we have the resolution
$0\to{\mathcal{O}}_{\mathbb{P}}(-12)\to{\mathcal{O}}_{\mathbb{P}}(-6)^{\oplus 2}\to{\mathcal{O}}_{\mathbb{P}}\to{\mathcal{O}}_{C}\to 0.$
By [Fle00] we have $\omega_{\mathbb{P}}={\mathcal{O}}_{{\mathbb{P}}}(-2-2-3-3)={\mathcal{O}}_{\mathbb{P}}(-10)$ and by Serre duality the subvariety $C$ of codimension two has canonical sheaf
@@
The above resolution computes this as $\omega_{C}\cong\mathcal{O}_{{\mathbb{P}}}(2){\raisebox{-0.1pt}{$|_{C}$}}$, which is invertible at all points of $C$, namely $C$ is Gorenstein. Finally, $h^{0}({\mathcal{L}})=h^{0}({\mathcal{O}}_{{\mathbb{P}}}(1))=0$.
@@ 5.E.3. $B(0)$
Computing the resolution of this ideal (e.g., taking the Eagon–Northcott complex or using MAGMA or Macaulay2, see the attached file), one sees that our condition on $g_{8}$ implies that it is a prime ideal, and computing the Hilbert series we see that it defines an irreducible curve $D_{1}$ with $p_{a}(D_{1})=1$. In particular $D_{1}$ is Gorenstein by [Cat82, Lemma 1.19]. An analogous computation for $C_{2}$ shows that $C$ contains $D_{1}\cup D_{2}$, where $D_{1}$ and $D_{2}$ are irreducible Gorenstein curves with arithmetic genus 1 and $C=D_{1}\cup D_{2}$ except possibly at the intersection of $C$ with $\{y_{1}=y_{2}=0\}$. This intersection consists only of the point $P:=[0,0,0,0,1,0,0,\xi]$, where $\alpha\xi^{3}=1$, and an explicit computation shows that $C$ has a node at $P$ since for a generic $g_{8}$ the two curves are smooth and transverse at $P$. Hence $C$ is reduced at all points, $C_{i}=D_{i}$, $i=1,2$, and $C=C_{1}\cup C_{2}$. Summing up, we have shown that $C$ is a reduced Gorenstein curve of genus two of type $B$. Hence $\omega_{C}$ is a line bundle.
@@ 5.E.5. $B(2)$
Data availability
During the work on this publication, no data sets were generated, used or analyzed. Thus, there is no need for a link to a data repository.
Declarations
Conflict of interest. On behalf of all authors, the corresponding author states that there is no conflict of interest.